# Gravity as a topological gauge theory

### v1

@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
$S[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi^{cd}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\phi^{ab}\wedge\phi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}-\dfrac{4}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}$ (27)
where the fact that $\alpha^{~{}\,a}_{~{}~{}\,\,\mu}\beta^{\,\mu}_{~{}\,\,b}\neq\dfrac{k}{2}\delta^{a}_{b}$ turns out to be an obstruction for $\varepsilon_{abcd}\phi^{ab}\wedge\phi^{cd}$ to be a constant volume form.
@@ 4 Conclusion
We have shown that a large class of models for gravitation, including the standard theory of general relativity, can be solely derived from topological considerations. This has been achieved by considering an action for gravitation as a linear combination of the Euler and Pontrjagin numbers of a manifold $\mathscr{M}$ with specific underlying $G/H$ Cartan geometry. Since the action is expressed as a linear combination of characteristic numbers constructed with various parts of the curvature of a Cartan connection, the action is invariant under the choice of such a Cartan connection up to the integral over $\mathscr{M}$ of an exact term. Two classes of examples have been studied from which Einstein’s equations (with positive or negative bare cosmological constant $\Lambda_{0}$) can be obtained upon making a restrictive choice (no Pontrjagin number). The first example corresponds to the case of a $\mathfrak{g}=\left\{\begin{array}[]{ll}\mathfrak{so}(4,1)\text{ for }\Lambda_{0}>0\\ \mathfrak{so}(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$-valued Cartan connection $\varpi$ defined on the $H=SO(3,1)$-principal-bundle $\mathscr{P}$. The computed action gives the Holst action together with the Euler, Pontrjagin and Nieh-Yan topological terms, see (15). The coupling constants of these topological terms are entirely determined by the two parameters of the Holst action ($G$ and $\gamma$) and the bare cosmological constant $\Lambda_{0}$. The second example deals with the Möbius geometry. In this case a very peculiar choice of the secondary tetrad $\alpha$ within the $\mathfrak{g}=\mathfrak{so}(4,2)$-valued Cartan connection $\varpi$ defined on the $H=CO(4,2)\ltimes\mathbb{R}^{3,1}$-principal-bundle $\mathscr{P}$ yields, either the same results as in the previous class of examples if one drops out dilations, or the Einstein’s equations with an additional source term for the curvature depending on a particular variation of spin density $-\dfrac{\pi G}{6}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;kc}$ when coupled to matter. In these two classes of examples treated in the paper, the action as linear combination of the Euler and Pontrjagin numbers contains the Holst action showing that the Holst action may have a topological origin. This opens the possibility to figure out a topological formulation of LQG.

### v3

@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
$S[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi^{cd}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\phi^{ab}\wedge\phi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}\bigr{)}-\dfrac{4r}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2rd\lambda\wedge d\lambda\Bigr{)}\biggr{)}$ (37)
where the fact that $\alpha^{a}{}_{\mu}\beta^{\,\mu}_{~{}\,\,b}\neq\dfrac{k}{2}\delta^{a}_{b}$ turns out to be an obstruction for $\varepsilon_{abcd}\phi^{ab}\wedge\phi^{cd}$ to be a constant volume form.
@@ 4 Conclusion
We have shown that a large class of models for gravitation, including the standard theory of general relativity, can be solely derived from topological considerations. This has been achieved by considering an action for gravitation as a linear combination of the Euler and Pontrjagin numbers of a manifold $\mathscr{M}$ with specific underlying $G/H$ Cartan geometry. Since the action is expressed as a linear combination of characteristic numbers constructed with various parts of the curvature of a Cartan connection, the action is invariant under the choice of such a Cartan connection up to the integral over $\mathscr{M}$ of an exact term.
Two classes of examples have been studied from which Einstein’s equations (with positive or negative bare cosmological constant $\Lambda_{0}$) can be obtained upon making a restrictive choice (no Pontrjagin numbers).
The first example corresponds to the case of a $\mathfrak{g}=\left\{\begin{array}[]{ll}\mathfrak{so}(4,1)\text{ for }\Lambda_{0}>0\\ \mathfrak{so}(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$-valued Cartan connection $\varpi$ defined on the $H=SO(3,1)$-principal-bundle $\mathscr{P}$. The computed action gives the Holst action together with the Euler, Pontrjagin and Nieh-Yan topological terms as well as a bare cosmological constant term, see (16). The coupling constants of these topological terms are entirely determined by the three parameters entering the Holst action ($G$, $\gamma$, $r$) and the bare cosmological constant $\Lambda_{0}$.
The second example deals with the Möbius geometry. In this case a very peculiar choice of the "secondary" tetrad $\alpha$ within the $\mathfrak{g}=\mathfrak{so}(4,2)$-valued Cartan connection $\varpi$ defined on the $H=CO(4,2)\ltimes\mathbb{R}^{3,1}$-principal bundle $\mathscr{P}$ yields, either the same results as in the previous class of examples if one drops out dilations, or the Einstein’s equations (modified by the Holst and Nieh-Yan terms) with an additional source term (due to the dilational part of the algebra) for the curvature depending on a particular variation of spin density $-\dfrac{2\pi G}{3}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;kc}$ and on torsion (if one does not assume $T=0$) when coupled to matter.
In these two classes of examples treated in the paper, the action as linear combination of the Euler and Pontrjagin numbers contains the Holst action showing that the Holst action may have a topological origin. This opens the possibility to figure out a topological formulation of LQG.
Acknowledgements
We thank our colleagues Thomas Schücker, Pietro Dona and Federico Piazza for helpful discussions.